Calculus II Spring 2002 MA114:004–006 Russell Brown

• Homework #2 §6.2 #27, 29, 31, 43, 49, 50, 56, 67, 69, 71, 83. §6.3 #1, 3, 15, 16, 17, 18, 35, 37, 59, 81, 82. To be discussed Tuesday, 15 January and Thursday, 17 January.

• Quiz #1. Thursday, 17 January 2002. Topic: §6.1. Finding derivatives of inverse functions.

• Homework #3 §6.4 #21, 23, 25, 27, 55, 58, 73, 75, 77, 91, 92. Lecture on Friday, 18 January. Discuss on Tuesday, 22 January.

January 10, 2002